OpenStudy (anonymous):

What is the sum of the infinite geometric series? -3-3/2-3/4-3/8-3/16- Help please. @Easyaspi314

OpenStudy (anonymous):



OpenStudy (anonymous):

I must leave now. Will be back later on in the afternnon or this evening.

OpenStudy (cwrw238):

the common ratio is (-3/2) / -3 = ½ sum to infinity of a GP = a1 / ( 1 - r) where a1 = first term , and r = common ratio

OpenStudy (anonymous):

Would I get a fraction? And okay! thanks @Easyaspi314

OpenStudy (cwrw238):

well its